vapor pressure of water at 293k is 17.535 mm hg.calculate the vapor pressure of water at 293k when 25g of glucose is dissolved in 450g of water.
1
Expert's answer
2020-04-24T13:33:48-0400

(P0 - P1)/P0=n(glucose )/(n(glucose)+n(H2O))

P0=17.535 mm Hg

n(glucose) = 25/180=0.1389 mol

n(H2O)=450/18=25 mol


P1=P0 - (n(glucose)*P0)/(n(glucose)+n(H2O)) =17.535 - (0.1389*17.535)/(0.1389+25)=17.438 mm Hg=2,325 kPa

Answer: 17.438 mm Hg or 2,325 kPa